# Unitarity, Causality, and Solar System Bounds, Significantly Limit Using Gravitational Waves to Test General Relativity

### v3

@@ Appendix C Discussion on $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$
In much of the literature when it comes to EFT treatment of gravity, any curvature corrections containing four-derivatives i.e. $R^{2}$, $R_{\mu\nu}R^{\mu\nu}$, and $(\text{Riem})^{2}$ are completely negated. The reasoning for the latter being forgotten is because it can be put into the form of the Gauss-Bonnet term which is a total derivative in the action. The same cannot be said about $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$. Instead, as in Stelle:1977ry , if we write down an action of the form $S\supset-\int\sqrt{-g}\left(\alpha R_{\mu\nu}R^{\mu\nu}-\beta R^{2}-\gamma\kappa^{-2}R\right)$, then the coefficients are $\gamma=2$ corresponding to Einstein gravity, while $\alpha$ and $\beta$ can be related to massive bodies as $m_{2}=\sqrt{\gamma/(\alpha\kappa^{2})}$ and $m_{0}=\sqrt{\gamma/(2(3\beta-\alpha)\kappa^{2}}$. If we consider the specific case of a point-particle, then the contribution to Newton’s potential is given by

### v5

@@ Appendix C Discussion on $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$
In much of the literature when it comes to EFT treatment of gravity, any curvature corrections containing four-derivatives i.e. $R^{2}$, $R_{\mu\nu}R^{\mu\nu}$, and $(\text{Riem})^{2}$ are completely neglected. The reasoning for the latter being forgotten is because it can be put into the form of the Gauss-Bonnet term which is a total derivative in the action. The same cannot be said about $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$. Instead, as in Stelle:1977ry , if we write down an action of the form $S\supset-\uprightintop\displaylimits\sqrt{-g}\left(\alpha R_{\mu\nu}R^{\mu\nu}-\beta R^{2}-\gamma\kappa^{-2}R\right)$, then the coefficients are $\gamma=2$ corresponding to Einstein gravity, while $\alpha$ and $\beta$ generate spin-2 and spin-0 massive excitations with masses $m_{2}=\sqrt{\gamma/(\alpha\kappa^{2})}$ and $m_{0}=\sqrt{\gamma/(2(3\beta-\alpha)\kappa^{2}}$. If we consider the specific case of a point-particle, then the contribution to Newton’s potential is given by
@@
Appendix D Distribution Functions
The function $P_{e}(y)$ (with $y=m_{\text{min}}/f$) is given by
$P_{e}(y)={6-6\,y\,\exp(y)\,\Gamma[0,y]\over 2-y+y^{2}-y^{3}\,\exp(y)\,\Gamma[0,y]}$ (D1)
where $\Gamma$ is the incomplete gamma function $\Gamma[0,y]=\displaylimits{}_{y}^{\infty}dt\,e^{-t}/t$.